OpenStudy (anonymous):

If you flip three fair coins, what is the probability that you'll get a head on the first flip, a tail on the second flip, and another head on the third flip?

hartnn (hartnn):

what is the probability of getting a head on 1 flip ?

OpenStudy (anonymous):

\[\left(\frac{1}{2}\right)^3=\frac{1}{8} \]

OpenStudy (mathmale):

Robert: As before, please help other users find their own answers. Your tendency to do others' work for them and to hand out answers without their involvement is problematic; it does little or nothing to teach understanding of the subject matter. Thank you.